# Share Your Attention: Transformer Weight Sharing via Matrix-based Dictionary Learning

### v1

@@ Results
- The output (O) projection matters: Comparing the two setups, MASA-QKV (unshared O) outperforms MASA-QKVO at similar compression rates. For example, at $S=4$, MASA-QKV achieves 121.4 perplexity on WikiText vs. 133.6 for MASA-QKVO. Thus, compressing the output projection (O) introduces a bottleneck that harms language modeling more than compressing Q, K, V.
@@
64_ (12 vs MASA)256_ (12 vs MASA)256_ (24 vs MASA)363638384040424244444646484850505252545439.4041.7544.6346.3546.7548.73Hidden Dimension & DepthAccuracy (%)Vanilla Attention Vs MASA on CIFAR100Vanilla 12 layersMASA-QKVO (S=4)02244668810100.6000.3284.7704.3049.5207.476Number of Parameters (M)
Figure 2: Evaluation results of different ViT models trained from scratch on CIFAR100 train data, the blue solid plot represents the Top1-Accuracy of the vanilla attention models, the green solid plot represents the Top1-Accuracy of MASA, the dotted lines represent the parameter count of the full models respectivly.
As illustrated in Fig. 2, our proposed method consistently surpasses vanilla attention across all scales. In particular, we investigated three experimental configurations:
  • A compact architecture with LL=12 layers and reduced width (hidden and MLP dimensions).
  • A fixed-depth model (LL=12) with $4$ times larger width (hidden and MLP dimensions).
  • An increased-depth (LL=24 layers) with preserved width.
@@
In Table 3, the proposed method demonstrates consistent superiority over SVD-LLM across different model scales, ain by both average accuracy on a diverse downstream benchmarks and perplexity on language modeling tasks. For larger architectures such as LLama 3.1 8B, our approach enables up to $20\%$ compression of the attention weight matrices while preserving approximately $99\%$ of the original model’s accuracy, indicating minimal degradation in semantic and reasoning capabilities despite parameter reduction.

### v2

@@ Results
- The output (O) projection matters: Comparing the two setups, MASA-QKV (unshared O) outperforms MASA-QKVO at similar compression rates. For example, at $S=4$, MASA-QKV achieves 121.4 perplexity on LAMBADA vs. 133.6 for MASA-QKVO. Thus, compressing the output projection (O) introduces a bottleneck that harms language modeling more than compressing Q, K, V.
@@
- Computational Overhead: MASA-QKVO achieves 1240 tokens/sec vs. 1352 tokens/sec for Transformer-S (sequence length nn=256, batch 16, A100-40GB), a ${\sim}8.3\%$ drop due to atom-based weighting. FLOPs increase from $4n^{2}d+8d^{2}n$ to $4n^{2}d+8d^{2}(n+k)$, where, $d$ is hidden dimension, Given that $k\ll n$ is the number of atoms, the additional computational cost remains negligible.
@@
64_ (12 vs MASA)256_ (12 vs MASA)256_ (24 vs MASA)363638384040424244444646484850505252545439.4041.7544.6346.3546.7548.73Hidden Dimension & DepthAccuracy (%)Vanilla 12 layersMASA-QKVO (S=4)02244668810100.6000.3284.7704.3049.5207.476Number of Parameters (M)
Figure 2: Evaluation results of different ViT models trained on CIFAR100 trainset, the blue and green solid plots represent the Top1-Accuracy of the vanilla attention models and MASA respectively. The dotted lines represent the parameter count of the full models respectively.
As illustrated in Fig. 2, our proposed method consistently surpasses vanilla attention across all scales. In particular, we investigated three experimental configurations: a compact architecture with LL=12 layers and reduced width (both hidden and MLP dimensions); a fixed-depth model with (LL=12) with $4$ times larger width in both hidden and MLP dimensions; and an increased-depth variant with (LL=24 layers) layers while preserving the original width.
@@
In Table 3, our method demonstrates consistent superiority over SVD-LLM across different model scales and compression ratios, achieving higher average accuracy across diverse downstream benchmarks while maintaining better language modeling performance as measured by perplexity. For larger architectures such as LLama 3.1 8B, our approach enables up to $20\%$ compression of the attention weight matrices while preserving approximately $99\%$ of the original model’s accuracy, indicating minimal degradation in semantic and reasoning capabilities despite parameter reduction.
@@ Conclusion
  • ARC-Easy (Clark et al. 2018): A subset of the AI2 Reasoning Challenge dataset consisting of science exam questions for which the correct answer can be inferred from a single sentence or common knowledge, formatted as multiple-choice problems.
  • SciQ (Welbl, Liu, and Gardner 2017): Evaluates scientific question answering by requiring models to select the correct answer from four options, based on science curriculum content typically encountered in middle and high school.
  • RACE (Lai et al. 2017): Assesses reading comprehension through multiple-choice questions derived from English exams for middle and high school students in China, covering diverse topics and requiring deep reasoning over long passages.